Sally’s Silk Screening produces specialty T-shirts that are primarily sold at special events. She is trying to decide how many to produce for an upcoming event. During the event, Sally can sell T-shirts for $20 apiece. However, when the event ends, any unsold T-shirts are sold for $4 apiece. It costs Sally $8 to make a specialty T-shirt. Sally’s estimate of demand is the following:

DEMAND PROBABILITY
300 0.05
400 0.10
500 0.40
600 0.30
700 0.10
800 0.05


a. What is the service rate (or optimal fractile)? (Round your answer to 2 decimal places.)

Service rate:
Answer #1

Service rate = cost of shortage / cost of shortage + cost of excess

cost of shortage = sales price - cost = 20-8 =12

cost of excess = cost - salvage value =8-4 =4

service rate = 12 / 12+4 = 12/16 = 0.75